ABC CORP BALANCE SHEET December 31, 2010 $ 100 Cash Marketable Securities 0 Accounts Receivables 2,000 Inventory 140 Fixed Asof Goods Sold. Fixed Operating Costs, excluding Depreciation Expense, were equal to $1,200 for 2010. These Fixed Operating CoThe Cash Budget Jan 11 Nov 10 Dec 10 $2,000 $1,000 Sales $4,000 Cost of goods sold 1,400 2,800 Beginning Inventory 280 560 EnLoan Required X X Loan Repaid X X Loan balance 0 Securities Purchased X X Securities Sold X X Securities Balance 0
Answer #1

Answer:

The Cash Budget

Note No Nov 10 ($) Dec 10 ($) Jan 11 ($)
Sales 2,000 4,000 1,000
Cost of Goods Sold 1,400 2,800

=1,000 x 70%

= 700

Beginning Inventory 280 560
Ending Inventory 560 140
Purchases 1,680 2,380
Cash Collections 1

=1,000+1,000

=2,000

=2,000+1,000

=3,000

=500+2,000

=2,500

Collected in Month of Sale

=50% x 2,000

= 1,000

=50% x 4,000

= 2,000

=50% x 1,000

=500

Collected month after sale

= 50% x 2,000

= 1,000

=50% x 2,000

=1,000

=50% x 4,000

= 2,000

Other Inflow
Interest Income Payments 0 0 0
Total Inflows 2,000 3,000
Outflows
Payment for Purchases 2

1,680

2,380

Interest Payments 3 0 0 0
Overheads Payments 4 100 100 100
Fixed Asset Additions 5 0 0 0
Dividend Payments 6 0 15 0
Income Tax Payments 7 0 0 0
Total Outflows 1,780 2495
Inflow - Outflow 220 505
Beginning Cash

=10% x 2,000

=200

20
Desired ( Ending ) Cash 8

=10% x 4,000

=400

100
Cash produced over + or Under - Immediate Need =220+200-400 =20 =20+505-100 = 425
Loan Required 0 0
Loan Repaid 0 0
Loan Balance 0 0
Securities Purchased 0 0
Securities Sold 0 0
Securities Balance 0 0

Notes :

1. Cash Collection: It is given that one half of the sales are collected in month of sale and remainder in following month

2. All purchases made are paid in month of following purchase means in same month

3. It is given that Interest Payment in year 2010 be $ 40. Therefore due to lack of information we assumed that there is no interest cost in month of nov 10 and dec 10 and also there will be no loan balance in that month

4. As per given information overheads includes Fixed costs , variable overheads are not given . Therefore , we consider only fixed costs

5.No fixed Asset addition in month of Nov 10 and Dec 10 as there is no specific information given

6. It is given that dividend will be paid at end of each quarter. For December 2010 Dividend amount paid be 15 ( 60/4 )

7. It is given that due to losses in ABC Corp as per the tax code the losses carry forward , therefore there is no Income Tax Payments and no Income Tax

8. Desired Cash Balance is 10% of Next Months Sale. Therefore,

Desired Cash Balance in :

Nov 10 :4000 ( dec sale ) x 10% = 400

Dec 10 : 1,000 ( Jan 11 sale ) x 10% = 100

Beginning cash Balance for Nov 10 be the desired cash balance ending for the month of Oct 10 as there is no information for oct 10 month

Beginning cash balance in Nov 10 =   2000 ( Oct Sale ) x 10% = 200